# On the principle of linearized stability for quasilinear evolution equations in time-weighted spaces

### v1

@@ 1. Introduction
$\overline{\mathbb{B}}_{E_{\alpha}}(v_{*},2\varepsilon/{\sf e}_{\alpha,\beta}\big{)}\subset\overline{\mathbb{B}}_{E_{\beta}}(v_{*},2\varepsilon\big{)}\subset O_{\beta}\,.$ (2.3)
We define
$\widehat{A}(w):=A(w+v_{*})+(\partial A(v_{*})[\cdot])v_{*}+\partial f(v_{*})\,,\qquad w\in\widehat{O}_{\beta}\,,$
and
$\widehat{f}(w):=f(w+v_{*})+A(w+v_{*})v_{*}-\partial f(v_{*})w-(\partial A(v_{*})[w])v_{*}\,,\qquad w\in\widehat{O}_{\xi},$
where we set $\widehat{O}_{\theta}:=O_{\theta}-v_{*}$ for $\theta\in\{\beta,\xi\}$. The functions $\widehat{f}$ and $\widehat{A}$ will appear in the reformulation (2.16) below of (1.1). Since $(\partial A(v_{*})[\cdot])v_{*}-\partial f(v_{*})\in\mathcal{L}(E_{\xi},E_{0})$ with $\xi<1$, we may use a perturbation argument [4, I.Theorem 1.3.1] to infer from (1.2b) that
$\widehat{A}\in C^{1-}\big{(}\widehat{O}_{\beta},\mathcal{H}(E_{1},E_{0})\big{)}\,.$ (2.4)
Let $\rho\in(0,\alpha-\beta)$ be fixed and note that $\alpha\geq\alpha_{\rm crit}^{*}$ implies
$\mu:=\xi-\alpha\leq\frac{1+\gamma_{*}-\alpha}{q_{*}}\,.$ (2.5)
In fact, the inequality in (2.5) turns into a strict inequality if $\alpha>\alpha_{\rm crit}^{*}$, that is, if $\gamma_{*}\in(0,\gamma)$ by assumption. If $\gamma_{*}\in(0,\gamma)$, we may thus choose $\gamma_{0}\in(0,\gamma_{*})$ such that $\mu q_{*}<1+\gamma_{0}-\alpha$, while we set $\gamma_{0}:=\gamma_{*}$ if ${\gamma_{*}\in\{0,\gamma\}}$. We then have in any case that
$\mu q_{*}\leq 1+\gamma_{0}-\alpha\,.$ (2.6)
Further, fix an arbitrary $\omega\in(0,\omega_{0})$ with $\omega_{0}-\omega=:4\delta>0$. Noticing that $\widehat{A}(0)=\mathbb{A}$, we infer from (2.4) and [4, I.Proposition 1.4.2] that there are $\kappa\geq 1$ and $\ell>0$ such that (making $\varepsilon>0$ smaller, if necessary) [^1] Given $\kappa\geq 1$ and $\varpi>0$, the set $\mathcal{H}(E_{1},E_{0};\kappa,\varpi)\subset\mathcal{H}(E_{1},E_{0})$ consists of the bounded operators $A\in\mathcal{L}(E_{1},E_{0})$ with the property $\varpi-A\in\mathcal{L}(E_{1},E_{0})$ is an isomorphism and $\kappa^{-1}\big{(}|\lambda|\,\|x\|_{0}+\|x\|_{1}\big{)}\leq\|(\lambda-A)x\|_{0}\leq\kappa\big{(}|\lambda|\,\|x\|_{0}+\|x\|_{1}\big{)},\qquad x\in E_{1}\,,\quad\mathrm{Re}\,\lambda\geq\varpi\,.$
$\omega_{0}-\delta+\widehat{A}(w)\in\mathcal{H}(E_{1},E_{0};\kappa,\delta)\,,\qquad w\in\overline{\mathbb{B}}_{E_{\beta}}(0,2\varepsilon)\,,$ (2.7)
and
$\|\widehat{A}(w_{1})-\widehat{A}(w_{2})\|_{\mathcal{L}(E_{1},E_{0})}\leq\ell\|w_{1}-w_{2}\|_{\beta}\,,\qquad w_{1},w_{2}\in\overline{\mathbb{B}}_{E_{\beta}}(0,2\varepsilon)\,.$ (2.8)
For $T\in(0,\infty)$, we define
[…]
where $N>0$ will be chosen below. More precisely, given $w\in\mathcal{M}(T)$, we derive from (2.7)-(2.8) that
$\omega_{0}-\delta+\widehat{A}(w(t))\in\mathcal{H}(E_{1},E_{0};\kappa,\delta)\,,\qquad t\in[0,T]\,,$ (2.9a)
and
$\widehat{A}(w)\in C^{\rho}\big{(}[0,T],\mathcal{L}(E_{1},E_{0})\big{)}\quad\text{with}\quad\sup_{0\leq s<t\leq T}\frac{\|\widehat{A}(w(t))-\widehat{A}(w(s))\|_{\mathcal{L}(E_{1},E_{0})}}{(t-s)^{\rho}}\leq N\,.$ (2.9b)
In view of (2.9) we may apply the results of [4, II.Section 5] to the family ${\mathcal{A}:=\{\widehat{A}(w)\,:\,w\in\mathcal{M}(T)\}}$. Letting $c_{0}(\rho)>0$ be the constant from [4, II.Theorem 5.1.1] (which is independent of $N$), we choose $N>0$ such that $c_{0}(\rho)N^{1/\rho}=\delta$. Then, by [4, II.Theorem 5.1.1, II.Lemma 5.1.3] there exists for each $w\in\mathcal{M}(T)$ a unique evolution operator $U_{\widehat{A}(w)}$ for $\widehat{A}(w)$ satisfying
$\begin{split}\|U_{\widehat{A}(w)}(t,s)\|_{\mathcal{L}(E_{\theta})}+(t-s)^{\theta-\vartheta_{0}}\|U_{\widehat{A}(w)}(t,s)\|_{\mathcal{L}(E_{\vartheta},E_{\theta})}\leq\frac{M_{1}}{2}e^{-\nu(t-s)}\,,\qquad 0\leq s\leq t\leq T\,,\end{split}$ (2.10a)
with a constant $M_{1}>0$ (independent of $T>0$) and
$-\nu:=c_{0}(\rho)N^{1/\rho}-\omega_{0}+\delta+\delta=-\omega-\delta<-\omega<0\,.$
The estimate (2.10a) holds for $0\leq\vartheta_{0}\leq\vartheta\leq\theta\leq 1$ with $\vartheta_{0}<\vartheta$ if $10<\vartheta<\theta<$ 1, or, thanks to (1.3), for $(\vartheta,\theta)\in\{(\alpha,\xi),(\gamma,\alpha),(\gamma,\beta),(\gamma,\xi)\}$ with $\vartheta_{0}=\vartheta$. Since we use (2.10a) only for a finite number of pairs $(\vartheta,\theta)$, we may assume that the constant $M_{1}$ does not depend on $(\vartheta,\theta)$. Moreover, [4, II.Theorem 5.3.1] (with ${f=0}$ therein) implies that there exists $M_{2}>0$ (independent of $T>0$) with
$\|U_{\widehat{A}(w)}(t,0)-U_{\widehat{A}(w)}(s,0)\|_{\mathcal{L}(E_{\alpha},E_{\beta})}\leq M_{2}(t-s)^{\alpha-\beta}\,,\quad 0\leq s\leq t\leq T\,,$ (2.10b)
while we may assume due to [4, II. Equation (5.3.8)] that
$\|U_{\widehat{A}(w)}(t,s)-1\|_{\mathcal{L}(E_{\alpha},E_{\beta})}\leq M_{2}(t-s)^{\alpha-\beta}\,,\quad 0\leq s\leq t\leq T\,.$ (2.10c)
Recalling (2.6), we may set
$c_{0}:=\sup_{r>0}\left(r^{1+\gamma_{0}-\alpha-\mu q_{*}}e^{-\nu r}\right)<\infty\,,$
and, with $\mathsf{B}$ denoting the Beta function and noticing that $01-\mu q_{*}\geq\alpha-\gamma_{0}\geq\alpha-\gamma_{*}\geq\alpha-\gamma>$ 0,
$\mathsf{B}_{\theta}:=\mathsf{B}(\theta,1-\mu q_{*}),\qquad\theta>0\,.$
Since $q_{*}>1$ and $\mu q_{*}\leq 1+\gamma_{0}-\alpha$, we can choose ${L>0}$ and ${\varepsilon_{0}\in(0,\varepsilon)}$ such that
$\frac{\varepsilon_{0}M_{1}}{2}+c_{*}c_{0}M_{1}L^{q_{*}}\mathsf{B}_{\mu q_{*}}\leq\frac{\varepsilon}{{\sf e}_{\alpha,\beta}}\,,$ (2.11a)
$\frac{\varepsilon_{0}M_{1}}{2}+c_{*}c_{0}M_{1}L^{q_{*}}\mathsf{B}_{\mu(q_{*}-1)}\leq\frac{L}{2}\,,$ (2.11b)
$\big{[}\varepsilon_{0}M_{2}+c_{*}c_{0}M_{1}L^{q_{*}}\big{(}M_{2}\mathsf{B}_{\mu q_{*}}+\mathsf{B}_{\alpha-\beta+\mu q_{*}}\big{)}\big{]}\Big{(}\frac{4\varepsilon\ell}{N}\Big{)}^{\tfrac{\alpha-\beta-\rho}{\rho}}\leq\frac{N}{2\ell}\,,$ (2.11c)
$c_{*}M_{1}L^{q_{*}-1}\big{(}\mathsf{B}_{\mu q_{*}}+\mathsf{B}_{\mu(q_{*}-1)}\big{)}\Big{(}\sup_{r>0}r^{1+\gamma_{0}-\alpha-\mu q_{*}}e^{(\omega-\nu)r}\Big{)}\leq\frac{1}{2}\,.$ (2.11d)
Also note from (1.8d)-(1.8e) that (assuming without loss of generality that $2\varepsilon/{\sf e}_{\alpha,\beta}\leq r_{*}$)
$\|\widehat{f}(w)\|_{\gamma_{*}}\leq 2c_{*}\|w\|_{\xi}^{q_{*}}\,,\qquad w\in\widehat{O}_{\xi}\cap\overline{\mathbb{B}}_{E_{\alpha}}(0,2\varepsilon/{\sf e}_{\alpha,\beta})\,.$ (2.12)
We will also use frequently the observation that
$\sup_{t>0}\left(t^{a}\int_{0}^{1}(1-s)^{-b}\,e^{-\varpi t(1-s)}\,s^{-\mu q_{*}}\,\mathrm{d}s\right)$ $\leq\Big{(}\sup_{r>0}r^{a}e^{-\varpi r}\Big{)}\,\mathsf{B}_{1-a-b}\,,\qquad 0\leq a<1-b\,,\quad\varpi>0.$ (2.13)
(b) Global Existence and Uniform Estimates. Let now $v^{0}\in\overline{\mathbb{B}}_{E_{\alpha}}(v_{*},\varepsilon_{0}\big{)}$ be fixed. Then, the evolution problem (1.1) has a unique maximal strong solution $v(\cdot;v^{0})$ on $[0,t^{+}(v^{0}))$ according to Theorem 1.2 with regularity properties stated in (1.5). Moreover, invoking Corollary 2.1 and making $\varepsilon_{0}>0$ smaller, if necessary, we may assume that $\overline{\mathbb{B}}_{E_{\alpha}}(v_{*},\varepsilon_{0}\big{)}\subset V_{\alpha}$ and $t^{+}(v^{0})>1$ with
$\|v(t;v^{0})-v_{*}\|_{\alpha}\leq k_{0}\varepsilon_{0}\leq 2\varepsilon/{\sf e}_{\alpha,\beta}\,,\qquad t\in[0,1]\,,\quad v^{0}\in\overline{\mathbb{B}}_{E_{\alpha}}(v_{*},\varepsilon_{0}\big{)}\,.$ (2.14)
Moreover, Corollary 2.1 ensures that there exist $t_{0}\in(0,1)$ and a constant $k_{1}>0$ such that
$\|v(t;v^{0})-v(s;v^{0})\|_{\beta}\leq k_{1}(t-s)^{\alpha-\beta},\qquad 0\leq s\leq t\leq t_{0}\,,\quad v^{0}\in\overline{\mathbb{B}}_{E_{\alpha}}(v_{*},\varepsilon_{0}\big{)}\,.$ (2.15)
We set $u^{0}:=v^{0}-v_{*}\in\overline{\mathbb{B}}_{E_{\alpha}}(0,\varepsilon_{0}\big{)}$ and $u:=v(\cdot;v^{0})-v_{*}$. Then $u$ is a strong solution to the Cauchy problem
$u^{\prime}=\widehat{A}(u)u+\widehat{f}(u)\,,\quad t>0\,,\qquad u(0)=u^{0}\,.$ (2.16)
Moreover, since $\rho<\alpha-\beta$, we can make $t_{0}\in(0,1)$ smaller, if necessary, to infer from (2.15) that
$\|u(t)-u(s)\|_{\beta}=\|v(t;v^{0})-v(s;v^{0})\|_{\beta}\leq k_{1}(t-s)^{\alpha-\beta}\leq\frac{N}{\ell}(t-s)^{\rho}\,,\qquad 0\leq s\leq t\leq t_{0}\,,\quad v^{0}\in\overline{\mathbb{B}}_{E_{\alpha}}(v_{*},\varepsilon_{0}\big{)}\,.$
Recalling (2.14), we deduce $u|_{[0,t_{0}]}\in\mathcal{M}(t_{0})$. Moreover, since $u\in C_{\mu}\big{(}(0,t_{0}],E_{\xi}\big{)}$ by Theorem 1.2, we may assume (making again $t_{0}>0$ smaller, if necessary) that $\|u\|_{C_{\mu}((0,t_{0}],E_{\xi})}\leq L$. Consequently, it holds that
$t_{1}:=\sup\big{\{}t<t^{+}(v^{0})\,:\,u|_{[0,t]}\in\mathcal{M}(t)\,,\,\|u\|_{C_{\mu}((0,t],E_{\xi})}\leq L\big{\}}\geq t_{0}\,.$
Let $t\in(0,t_{1})$. Then, $\|u(s)\|_{\xi}\leq s^{-\mu}L$ for $s\in(0,t]$, and (2.10) holds for $\widehat{A}(u)$ and $T=t$ since ${u|_{[0,t]}\in\mathcal{M}(t)}$. Moreover, $u$ satisfies the variation-of-constants formula
$u(\tau)=U_{\widehat{A}(u)}(\tau,0)u^{0}+\int_{0}^{\tau}U_{\widehat{A}(u)}(\tau,s)\widehat{f}(u(s))\,\mathrm{d}s\,,\quad 0\leq\tau\leq t\,,$
according to (1.5) and (2.16). Therefore, we deduce from (2.10a) and (2.12) for $\theta\in\{\alpha,\,\xi\}$ that
$\|u(t)-U_{\widehat{A}(u)}(t,0)u^{0}\|_{\theta}$ $\leq\int_{0}^{t}\|U_{\widehat{A}(u)}(t,s)\|_{\mathcal{L}(E_{\gamma_{*}},E_{\theta})}\,\|\widehat{f}(u(s))\|_{\gamma_{*}}\,\mathrm{d}s$
$\leq c_{*}M_{1}L^{q_{*}}\int_{0}^{t}(t-s)^{\gamma_{0}-\theta}\,e^{-\nu(t-s)}\,s^{-\mu q_{*}}\,\mathrm{d}s$
$=c_{*}M_{1}L^{q_{*}}\,t^{1+\gamma_{0}-\theta-\mu q_{*}}\int_{0}^{1}(1-s)^{\gamma_{0}-\theta}\,e^{-\nu t(1-s)}\,s^{-\mu q_{*}}\,\mathrm{d}s\,.$ (2.17)
Choosing $\theta=\alpha$ in (2.17), we infer from (2.10a) and (2.13) that
$\|u(t)\|_{\alpha}$ $\leq\frac{M_{1}}{2}e^{-\nu t}\|u^{0}\|_{\alpha}+c_{*}M_{1}L^{q_{*}}\,t^{1+\gamma_{0}-\alpha-\mu q_{*}}\int_{0}^{1}(1-s)^{\gamma_{0}-\alpha}\,e^{-\nu t(1-s)}\,s^{-\mu q_{*}}\,\mathrm{d}s$
$\leq\frac{\varepsilon_{0}M_{1}}{2}+c_{*}M_{1}L^{q_{*}}\,\Big{(}\sup_{r>0}r^{1+\gamma_{0}-\alpha-\mu q_{*}}e^{-\nu r}\Big{)}\mathsf{B}_{\mu q_{*}}\,,$
and, recalling (2.11a), we thus have
$\|u(t)\|_{\alpha}\leq\frac{\varepsilon}{{\sf e}_{\alpha,\beta}}\,,\qquad t\in(0,t_{1})\,.$ (2.18)
Now, since
$\|U_{\widehat{A}(u)}(t,0)u^{0}\|_{\xi}$ $\leq\frac{M_{1}}{2}t^{-\mu}e^{-\nu t}\|u^{0}\|_{\alpha}\,,$
due to (2.10a), we obtain from (2.17) (with $\theta=\xi$), using (2.13), that
$t^{\mu}\|u(t)\|_{\xi}$ $\leq\frac{M_{1}}{2}e^{-\nu t}\|u^{0}\|_{\alpha}+c_{*}M_{1}L^{q_{*}}\,t^{1+\gamma_{0}+\mu-\xi-\mu q_{*}}\int_{0}^{1}(1-s)^{\gamma_{0}-\xi}\,e^{-\nu t(1-s)}\,s^{-\mu q_{*}}\,\mathrm{d}s$
$\leq\frac{\varepsilon_{0}M_{1}}{2}+c_{*}M_{1}L^{q_{*}}\,\Big{(}\sup_{r>0}r^{1+\gamma_{0}-\alpha-\mu q_{*}}e^{-\nu r}\Big{)}\mathsf{B}_{\mu(q_{*}-1)}\,.$
Therefore, (2.11b) implies that
$\|u\|_{C_{\mu}((0,t],E_{\xi})}\leq\frac{L}{2}\,,\qquad t\in(0,t_{1})\,.$ (2.19)
Next, let $0\leq s<t<t_{1}$. If $|t-s|^{\rho}\geq 4\varepsilon\ell/N$ , then (2.18) yields
$\|u(t)-u(s)\|_{\beta}\leq 2\varepsilon=\frac{2\varepsilon}{|t-s|^{\rho}}|t-s|^{\rho}\leq\frac{N}{2\ell}|t-s|^{\rho}\,,\qquad|t-s|^{\rho}\geq\frac{4\varepsilon\ell}{N}\,.$ (2.20)
Conversely, consider $|t-s|^{\rho}\leq 4\varepsilon\ell/N$ and note that
$\|u(t)-u(s)\|_{\beta}$ $\leq\|U_{\widehat{A}(u)}(t,0)u^{0}-U_{\widehat{A}(u)}(s,0)u^{0}\|_{\beta}$
$\quad+\int_{0}^{s}\|U_{\widehat{A}(u)}(t,\tau)-U_{\widehat{A}(u)}(s,\tau)\|_{\mathcal{L}(E_{\gamma_{*}},E_{\beta})}\,\|\widehat{f}(u(\tau))\|_{\gamma_{*}}\,\mathrm{d}\tau$
$\quad+\int_{s}^{t}\|U_{\widehat{A}(u)}(t,\tau)\|_{\mathcal{L}(E_{\gamma_{*}},E_{\beta})}\,\|\widehat{f}(u(\tau))\|_{\gamma_{*}}\,\mathrm{d}\tau=:I_{1}+I_{2}+I_{3}\,.$ (2.21a)
From (2.10b) it follows that
$I_{1}\leq M_{2}\|u^{0}\|_{\alpha}\,(t-s)^{\alpha-\beta}\leq\varepsilon_{0}M_{2}\Big{(}\frac{4\varepsilon\ell}{N}\Big{)}^{\tfrac{\alpha-\beta-\rho}{\rho}}(t-s)^{\rho}\,,$ (2.21b)
while (2.10a), (2.10c), and (2.12) entail
$I_{2}$ $\leq\int_{0}^{s}\|U_{\widehat{A}(u)}(t,s)-1\|_{\mathcal{L}(E_{\alpha},E_{\beta})}\,\|U_{\widehat{A}(u)}(s,\tau)\|_{\mathcal{L}(E_{\gamma_{*}},E_{\alpha})}\|f(u(\tau))\|_{\gamma_{*}}\,\mathrm{d}\tau$
$\leq c_{*}M_{1}M_{2}L^{q_{*}}(t-s)^{\alpha-\beta}\int_{0}^{s}(s-\tau)^{\gamma_{0}-\alpha}e^{-\nu(s-\tau)}\tau^{-\mu q_{*}}\,\mathrm{d}\tau$
$=c_{*}M_{1}M_{2}L^{q_{*}}(t-s)^{\alpha-\beta}s^{1+\gamma_{0}-\alpha-\mu q_{*}}\int_{0}^{1}(1-\tau)^{\gamma_{0}-\alpha}e^{-\nu s(1-\tau)}\tau^{-\mu q_{*}}\,\mathrm{d}\tau$
and hence, using (2.13),
$I_{2}$ $\leq c_{*}c_{0}M_{1}M_{2}L^{q_{*}}\mathsf{B}_{\mu q_{*}}\Big{(}\frac{4\varepsilon\ell}{N}\Big{)}^{\tfrac{\alpha-\beta-\rho}{\rho}}(t-s)^{\rho}\,.$ (2.21c)
Finally, from (2.10a) and (2.12) we get
$I_{3}$ $\leq c_{*}M_{1}L^{q_{*}}\int_{s}^{t}(t-\tau)^{\gamma_{0}-\beta}e^{-\nu(t-\tau)}\tau^{-\mu q_{*}}\,\mathrm{d}\tau$
$\leq c_{*}M_{1}L^{q_{*}}(t-s)^{\alpha-\beta}(t-s)^{1+\gamma_{0}-\alpha-\mu q_{*}}\int_{0}^{1}(1-\tau)^{\gamma_{0}-\beta}e^{-\nu(t-s)(1-\tau)}\tau^{-\mu q_{*}}\,\mathrm{d}\tau$
and therefore, invoking again (2.13),
$I_{3}$ $\leq c_{*}c_{0}M_{1}L^{q_{*}}\mathsf{B}_{\alpha-\beta+\mu q_{*}}\Big{(}\frac{4\varepsilon\ell}{N}\Big{)}^{\tfrac{\alpha-\beta-\rho}{\rho}}(t-s)^{\rho}\,.$ (2.21d)
Consequently, gathering (2.20)-(2.21) and then using (2.11c), we deduce that
$\|u(t)-u(s)\|_{\beta}\leq\frac{N}{2\ell}(t-s)^{\rho},\qquad 0\leq s\leq t<t_{1}\,.$ (2.22)
Summarizing, we obtain from the estimates (2.18), (2.19), and (2.22) that $t_{1}=t^{+}(v^{0})$. In particular, the solution ${v(\cdot;v^{0}):[0,t^{+}(v^{0}))\to E_{\beta}}$ is uniformly continuous. Moreover, recalling (1.2d) and (2.18), and noticing from (2.19) that
$\|v(t;v^{0})\|_{\xi}\leq\|v_{*}\|_{\xi}+Lt^{-\mu}\,,\qquad t\in(0,t^{+}(v^{0}))\,,$
we also have
$\underset{t\nearrow t^{+}(v^{0})}{\limsup}\|f(v(t;v^{0}))\|_{0}<\infty$
while (2.3) ensures that
$\liminf_{t\nearrow t^{+}(v^{0})}\mathrm{dist}_{E_{\beta}}\big{(}v(t;v^{0}),\partial O_{\beta}\big{)}>0\,.$
Consequently, Theorem 1.2 implies that $t^{+}(v^{0})=\infty$. In particular, the evolution system $U_{\widehat{A}(u)}(t,s)$ is defined for all $0\leq s\leq t<\infty$ and satisfies (2.10a) for all $0\leq s\leq t<\infty$.
(c) Stability. Set
$z(t):=\sup_{\tau\in(0,t]}\big{(}\|u(\tau)\|_{\alpha}+\tau^{\mu}\|u(\tau)\|_{\xi}\big{)}e^{\omega\tau}\,,\qquad t>0\,,$
and note from (2.19) that
$\|u(\tau)\|_{\xi}^{q_{*}}\leq\tau^{-\mu q_{*}}e^{-\omega\tau}L^{q_{*}-1}z(t)\,,\qquad 0<\tau\leq t\,.$
Using the latter along with (2.10a) and (2.12), it follows analogously as before [^2] The exponential $e^{-\nu r}$ has now to be replaced everywhere by $e^{(\omega-\nu)r}$. when deriving (2.18) and (2.19) that for $0<\tau\leq t$ we have
$\big{(}\|u(\tau)\|_{\alpha}+\tau^{\mu}\|u(\tau)\|_{\xi}\big{)}e^{\omega\tau}$
$\quad\leq\frac{c_{1}}{2}\|u^{0}\|_{\alpha}+c_{*}M_{1}L^{q_{*}-1}\big{(}\mathsf{B}_{\mu q_{*}}+\mathsf{B}_{\mu(q_{*}-1)}\big{)}\Big{(}\sup_{r>0}r^{1+\gamma_{0}-\alpha-\mu q_{*}}e^{(\omega-\nu)r}\Big{)}z(t)\,,$
where $c_{1}:=2M_{1}$. Invoking (2.11d), we thus have
$\big{(}\|u(t)\|_{\alpha}+t^{\mu}\|u(t)\|_{\xi}\big{)}e^{\omega t}\leq z(t)\leq c_{1}\|u^{0}\|_{\alpha}\,,\qquad t>0\,,$
and, since $v(t;v^{0})=u(t)+v_{*}$ and $v^{0}=u^{0}+v_{*}$, the proof of Theorem 1.3 is complete.∎
@@
The proof of Theorem 1.6 follows along the lines of the proof of Theorem 1.3. We thus only sketch it.
@@
Let the assumptions of Theorem 1.6 hold. To start with, we note that since by assumption $\alpha\geq\alpha_{\rm crit}^{*}$, we have
@@
Let ${R\in(0,r_{*})}$, see (1.15d), be chosen such that $\overline{\mathbb{B}}_{E_{\alpha}}(v_{*},R)\subset O_{\alpha}$ and fix
$\omega_{0}>\zeta>\omega>0\,.$
@@
$c_{0}$ $:=1+\big{(}{\sf B}_{\mu q_{*}}+{\sf B}_{\mu(q_{*}-1)}\big{)}\Big{(}\sup_{r>0}r^{1+\gamma_{0}-\alpha-\mu q_{*}}e^{(\omega-\zeta)r}\Big{)}\,,$
with ${\sf B}_{\theta}:={\sf B}(\theta,1-\mu q_{*})$ for $\theta>0$, is well-defined. Assumption (1.15e) together with [4, II.Lemma 5.1.3] ensures the existence of a constant ${M\geq 1}$ such that the strongly continuous analytic semigroup $(e^{t\mathbb{A}})_{t\geq 0}$ generated by $\mathbb{A}$ satisfies
$\|e^{t\mathbb{A}}\|_{\mathcal{L}(E_{\theta})}+t^{\theta-\vartheta_{0}}\|e^{t\mathbb{A}}\|_{\mathcal{L}(E_{\vartheta},E_{\theta})}\leq\frac{M}{4c_{0}}e^{-\zeta t}\,,\qquad t>0\,,$ (3.1)
for $0\leq\vartheta_{0}\leq\vartheta\leq\theta\leq 1$ with $\vartheta_{0}<\vartheta$ if $10<\vartheta<\theta<$ 1 or, thanks to (1.12), for $(\vartheta,\theta)\in\{(\alpha,\xi),(\gamma,\alpha),(\gamma,\xi)\}$ with $\vartheta_{0}=\vartheta$. We may then fix $L\in(0,R)$ such that
@@
where $c_{*}>0$ stems from (1.15d). We now set
@@
see (1.15d). The estimates (2.13), (3.1), and (3.4) lead to

### v2

@@ 1. Introduction
for each $v^{0}\in\overline{\mathbb{B}}_{E_{\alpha}}(v_{*},\varepsilon_{0})$, which proves Theorem 1.3.
@@
This section is devoted to the proof of Theorem 1.6. Under the assumption (a) of this theorem, the statement follows along the same lines as the proof of Theorem 1.3. As a preliminary result, we therefore establish in Proposition 3.1 the counterpart of Proposition 2.1 and Corollary 2.2 in the semilinear case.
Proposition 3.1.
Assume (1.11) with $\xi<1$ and (1.12). Then, given $\overline{v}\in O_{\alpha}$, there exist $r=r(\overline{v})>0$} and $T=T(\overline{v})>0$ such that, for each $v^{0}\in\bar{\mathbb{B}}_{E_{\alpha}}(\overline{v},r)\subset O_{\alpha}$, the Cauchy problem (1.10) possesses a unique solution
$v(\cdot;v^{0})$ $\in C\big((0,T],E_{1}\big)\cap C^{1}\big((0,T],E_{0}\big)\cap C\big([0,T],O_{\alpha}\big)\cap C_{\xi-\alpha}\big((0,T],E_{\xi}\big)\,.$ (3.1)
Moreover, there exists a constant $c=c(\overline{v})>0$ such that, for $v^{0},\,v^{1}\in\bar{\mathbb{B}}_{E_{\alpha}}(\overline{v},r)$, it holds that
$\|v(\cdot;v^{0})-v(\cdot;v^{1})\|_{C([0,T],E_{\alpha})}+\|v(\cdot;v^{0})-v(\cdot;v^{1})\|_{C_{\xi-\alpha}((0,T],E_{\xi})}\leq c\|v^{0}-v^{1}\|_{\alpha}.$ (3.2)
Fixing an arbitrary $\zeta\in(\xi,1)$ it additionally holds, for $v^{0},\,v^{1}\in\bar{\mathbb{B}}_{E_{\alpha}}(\overline{v},r)$, that
$\|v(\cdot;v^{0})-v(\cdot;v^{1})\|_{C_{\zeta}((0,T],E_{\zeta})}\leq c\|v^{0}-v^{1}\|_{\alpha}.$ (3.3)
Proof.
The claims, up to the estimate (3.3), are established in [14, Proposition 3.1] for $\alpha_{\rm crit}<\alpha$ and in [12, Proposition 3.1] for $\alpha_{\rm crit}=\alpha$, with the observation that in the noncritical case $\alpha_{\rm crit}<\alpha$ the regularity property (3.1) and the estimate (3.2) in $C_{\xi-\alpha}((0,T],E_{\xi})$ are obtained by also exploiting (1.12) (see the proof of (3.3) below).
In order to prove (3.3), we infer from (1.11e) and (3.2), that there exists a constant $c=c(\bar{v})$ such that for all $v^{0},\,v^{1}\in\bar{\mathbb{B}}_{E_{\alpha}}(\overline{v},r)$ we have
$\|f(v(t;v^{0}))-f(v(t;v^{1}))\|_{0}\leq\|f(v(t;v^{0}))-f(v(t;v^{1}))\|_{\gamma}\leq c_{2}t^{-q(\xi-\alpha)}\|v^{0}-v^{1}\|_{\alpha}\,,\quad 0<t\leq T\,.$ (3.4)
Since solutions to (1.10) satisfy the variation of constant formula
$v(t;v^{0})=e^{tA}v^{0}+\int_{0}^{t}e^{(t-s)A}f(v(s;v^{0}))\,\mathrm{d}s\,,\qquad t\in[0,T]\,$,
using (3.4) together with the estimate $\|e^{(t-s)A}\|_{\mathcal{L}(E_{0},E_{\zeta})}\leq C(t-s)^{-\zeta}$ for $0\leq s<t\leq T$, we get for $t\in(0,T]$ and $v^{0}\in\bar{\mathbb{B}}_{E_{\alpha}}(\overline{v},r)$ that
$t^{\zeta}\|v(t;v^{0})-v(t;v^{1})\|_{\zeta}$ $\leq t^{\zeta}\|e^{tA}\|_{\mathcal{L}(E_{0},E_{\zeta})}\,\|v^{0}-v^{1}\|_{0}$
$\qquad+t^{\zeta}\int_{0}^{t}\big\|e^{(t-s)A}\big\|_{\mathcal{L}(E_{0},E_{\zeta})}\,\|f(v(s;v^{0}))-f(v(s;v^{1}))\|_{0}\,\mathrm{d}s$
$\leq c\|v^{0}-v^{1}\|_{\alpha}+ct^{\zeta}\int_{0}^{t}(t-s)^{-\zeta}s^{-q(\xi-\alpha)}\,\mathrm{d}s\,\|v^{0}-v^{1}\|_{\alpha}$
$=c\big[1+t^{1-q(\xi-\alpha)}\mathsf{B}(1-\zeta,1-q(\xi-\alpha))\big]\|v^{0}-v^{1}\|_{\alpha}\,.$
Above, $\mathsf{B}$ denotes again the Beta function and we used the inequality $1-q(\xi-\alpha)>0$ which is a consequence of (1.11c)-(1.11d). ∎
We conclude this section with the proof of Theorem 1.6.
@@
If the assumption (a) holds, then the proof follows from Proposition 3.1, using a line of reasoning similar to that in the proof of Theorem 1.3.
We now prove Theorem 1.6, assuming (b). Since in this case $\alpha\geq\alpha_{\rm crit}^{*}$, we have
@@
Let $R\in(0,r_{*})$}, see (1.16b), be chosen such that $\overline{\mathbb{B}}_{E_{\alpha}}(v_{*},R)\subset O_{\alpha}$ and fix
$\omega_{0}>\bar{\omega}>\omega>0\,.$
@@
$c_{0}$ $:=1+\big({\sf B}_{\mu q_{*}}+{\sf B}_{\mu(q_{*}-1)}\big)\Big(\sup_{r>0}r^{1+\gamma_{0}-\alpha-\mu q_{*}}e^{(\omega-\bar{\omega})r}\Big)$
with ${\sf B}_{\theta}:={\sf B}(\theta,1-\mu q_{*})$ for $\theta>0$ ($\mathsf{B}$ denotes again the Beta function), is well-defined. Assumption (1.15c) together with [4, II.Lemma 5.1.3] ensure the existence of a constant $M\geq 1$} such that the strongly continuous analytic semigroup $(e^{t\mathbb{A}})_{t\geq 0}$ generated by $\mathbb{A}$ satisfies
$\|e^{t\mathbb{A}}\|_{\mathcal{L}(E_{\theta})}+t^{\theta-\vartheta_{0}}\|e^{t\mathbb{A}}\|_{\mathcal{L}(E_{\vartheta},E_{\theta})}\leq\frac{M}{4c_{0}}e^{-\bar{\omega}t}\,,\qquad t>0\,$, (3.5)
for $0\leq\vartheta_{0}\leq\vartheta\leq\theta\leq 1$ with $\vartheta_{0}<\vartheta$ if $0<\vartheta<\theta<1$ or, thanks to (1.12), for $(\vartheta,\theta)\in\{(\alpha,\xi),(\gamma,\alpha),(\gamma,\xi)\}$ with $\vartheta_{0}=\vartheta$. We may then fix $L\in(0,R)$ such that
@@
where $c_{*}>0$ stems from (1.16b). We now set
@@
see (1.16b). In view of the estimate
$\sup_{t>0}\left(t^{a}\int_{0}^{1}(1-s)^{-b}\,e^{-\eta t(1-s)}\,s^{-\mu q_{*}}\,\mathrm{d}s\right)$ $\leq\Big(\sup_{r>0}r^{a}e^{-\eta r}\Big)\,\mathsf{B}_{1-a-b}\,,\qquad 0\leq a<1-b\,,\quad\eta>0,$ (3.9)
we deduce from (3.5) and (3.8) that